This is my second assign.
Please help me solve this. Thank you again in advance
[5(2a-b)-3][5(2a-b)+3]

Found 4 solutions by MathLover1, greenestamps, MathTherapy, Alan3354:
Answer by MathLover1(20850)
Answer by greenestamps(13209)



If you look at the expression carefully, there is a path to the answer with much less work than shown by the other tutor.



This is in the form (A+B)(A-B) = A^2-B^2. So




Answer by MathTherapy(10557)


This is my second assign.
Please help me solve this. Thank you again in advance
[5(2a-b)-3][5(2a-b)+3]
[5(2a - b) - 3][5(2a - b) + 3] =  


Answer by Alan3354(69443)

Equations are solved, eg, "Solve for x"
This is not an equation.
I can tell - there is no equal sign.